An investment of $500 at 8% per annum compounded annually grows to A$ in n years
a) solve this equation for n
b) so expressing n as a log function of A
calculate the value of n for each value of A 1) A = 1250 2) A = 350
c) graph the function in part a for 0 d) state the domain and range
Can you please help me out ? Thanks so much in advance:)

Answer by stanbon(75887)

An investment of $500 at 8% per annum compounded annually grows to A$ in n years
Equation:: A = 500(1+0.08)^n
---------------------------------------
a) solve this equation for n
1.08^n = A/500
---------------------
b) so expressing n as a log function of A
n*log(1.08) = log[A/500]
n = [log(A)-log(500)]/log(1.08)
---------------------------------------------
calculate the value of n for each value of A
1) A = 1250
n = [log(1250)-log(500)]/log(1.08) = 11.91
-------------------------------------------
2) A = 350
n = [log(350)-log(500)]/log(1.08) = -4.6345
=================================================

c) graph the function in part a for 0
Note: 0 is not a variable of that equation.
----------------------------------------
d) state the domain and range
Domain: n is integer >= zero
Range: A is a Real Numbers >= 500
=================
Cheers,
Stan H.
-----------------